OpenStudy (anonymous):

find an equation of the plane with x-intercept (a,0,0), y-intercept (0,b,0), z-intercept (0,0,c) where a,b,c are positive real numbers.

OpenStudy (anonymous):

Im not sure where to start

OpenStudy (anonymous):

equation of a plane is Ax+By+Cz+D=0

OpenStudy (anonymous):

u have three points that satisfies this equation (x,y,z)=(a,0,0),(0,b,0),(0,0,c)

OpenStudy (anonymous):

for example A*a+B*0+C*0+D=0 then A=-D/a

OpenStudy (anonymous):

try to find B and C

OpenStudy (anonymous):

so B = -D/b ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(-D/a)x + -D/b(y) + (-D/c)z = 0

OpenStudy (anonymous):

emmm (-D/a)x + -D/b(y) + (-D/c)z+D=0

OpenStudy (anonymous):

now eliminate D

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

you have \[\frac{x}{a}D+\frac{y}{b}D+\frac{z}{c}D-D=0\]

OpenStudy (anonymous):

divide it by D

OpenStudy (anonymous):

what is result?

OpenStudy (anonymous):

z/a + y/b + z/c = 1

OpenStudy (anonymous):

almost done

OpenStudy (anonymous):

isnt that the answer? I was not given values for a,b, or c

OpenStudy (anonymous):

thats the answer a b and c are fixed constants

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw